OpenStudy (anonymous):

If m varies directly as the square root of y, inversely as p^2, and directly as n, what happens to m when y is quadrupled, p is tripled, and n is multiplied by 5?

OpenStudy (anonymous):

Answer choices:

OpenStudy (anonymous):

so... m = k y^0.5 p^-2 n? do this: m'=k (4y)^0.5 (3p)^-2 5n

OpenStudy (anonymous):

@RichGrl20 does that help?

OpenStudy (anonymous):

@Peter14 : Yes, it does! Thanks. :D

OpenStudy (anonymous):

glad to help. anything else?

OpenStudy (anonymous):

@Peter14 : Um... I'm not too sure how to figure out what happens to m. I know that each equation models what is said in the description, but I don't know how to use those to find out what change is applied to m, if any.

OpenStudy (anonymous):

apply the exponents to the numbers I put in

OpenStudy (anonymous):

just ignore k

OpenStudy (anonymous):

here, there's an easier way. go back to the first equation I made, except leave out the k

OpenStudy (anonymous):

m=y^0.5 p^-2 n plug in a 1 for each of those numbers. then go back and change each number as specified in the description of the problem at the top

OpenStudy (anonymous):

@RichGrl20 does that help?

OpenStudy (anonymous):

I may have done it wrong, but I'm not getting any of the answer choices...

OpenStudy (anonymous):

set it up for me

OpenStudy (anonymous):

@RichGrl20 it is one of the answer choices, do it here so I can see where you made a mistake

OpenStudy (anonymous):

nevermind, I have to go to bed now because it's 10:35 PM in China.